THEORETICAL AND EXPERIMENTAL STUDY OF THE ROVIBRATIONAL SPECTRA OF CO_{2}-(para-H_{2})-He TRIMERS

\underline{HUI LI}, Institute of Theoretical Chemistry, State Key Lab. of Theoretical & Computational Chemistry, Jilin Univ., 2519 Jiefang Rd, Changchun 130023, P.R.China; Chemistry Dept., Univ. of Waterloo, Waterloo, Ontario N2L 3G1, Canada; ROBERT J. LE ROY, PIERRE-NICHOLAS ROY, Chemistry Dept., Univ. of Waterloo, Waterloo, Ontario N2L 3G1, Canada; A. R. W. McKELLAR, Steacie Institute for Molecular Sciences, NRCC, Ottawa, Ontario K1A OR6, Canada.

Clusters of p-H_{2} had been predicted to exhibit superfluid behavior twenty years ago,^{a} but direct observation of this phenomenon was elusive until our recent work^{b} combining experimental measurments and theoretical simulations of the non-classical rotational inertia and superfluid response of p-H_{2} clusters doped with CO_{2}. However, the size-dependent superfluid response of those clusters reached a maxmum at N = 12, and the clusters become frozen at larger N. It is therefore interesting to examine the effect of adding helium atoms to a pure p-H_{2} shell around a CO_{2} chromophore, and to investigate their effect on the superfluid response of CO_{2}-(p-H_{2})_{N} for N >12. This will help us understand the role of helium as a ‘second solvent’ species, and help explain experiments that had been presented as evidence of superfluidity of doped hydrogen clusters embedded in helium nanodroplets.^{c} Exact quantum calculation of infrared and microwave spectra for dopant molecules attached to two pure He atoms or two pure p-H_{2} molecules, respectively, have been reported by Wang and co-workers and by Li et al.^{d} To date, however, no such calculations have been reported for mixed p-H_{2}/He solvent species. The present paper therefore extends such work to the case of one p-H_{2} and one He atom attached to one CO_{2} dopant molecule. Three-dimensional p-H_{2} and helium densities in the body-fixed frame are presented, and compared with those for the analogous CO_{2}–(He)_{2} and CO_{2}–(p-H_{2})_{2} trimers.

a P. Sindzingre, D.M. Ceperley, and M.L. Klein, Phys. Rev. Lett. 67, 1871 (1991).

b H. Li, R.J. Le Roy, P.-N. Roy, and A.R.W. McKellar, Phys. Rev. Lett. 105, 133401 (2010)

c S. Grebenev, B. Sartakov, J.P. Toennies, and A.F. Vilesov Science 289, 1532 (2000).

d X.-G. Wang, T. Carrington Jr., J. Tang and A.R.W. McKellar, Can. J. Phys. 87, 417 (2009); J. Phys. Chem. A 113, 13331 (2009); Can. J. Phys. 88, 779 (2010); H. Li, P.-N. Roy and R.J. Le Roy, J. Phys. Chem. (2011, submitted).